# Hybrid User Interfaces: Past, Present, and Future of Complementary Cross-Device Interaction in Mixed Reality

### v1

@@ 4.1 Taxonomy Dimensions
The relationship category denotes the relation between users within one or multiple systems. The vast majority of our corpus describes single-user systems 72: One user interacts with multiple complementary interaction components (see Fig. 4 1 user). However, we identified several collaborative systems. We abstracted these into multi-user with individual interaction components 7: Each collaborator has their own set of interaction components [130, 121] (see Fig. 4 n-ind.); and multi-user with shared interaction component 12: A display is shared between collaborators as a public space [101, 119, 85, 18] (see Fig. 4 n-shared).
@@
We examined the range of interaction across components as defined in the cross-device taxonomy [16]. Since the MR interaction component is commonly worn on the user’s head (e.g., AR HWD) or close to their body (e.g., handheld AR), we use this dimension to describe the scale of the 2D interaction component relative to the user. In addition, the relationship can be an indicator of the range between interaction components (e.g., multi-user can indicate social or public range). We differentiate between near 6 (i.e., interaction component is close to the user’s body, e.g., smartwatch [54, 145, 41]), personal 65 (i.e., interaction component is in the personal space, e.g., smartphone or tablet [124, 67, 87, 85]), social 23 (i.e., interaction component is in a social space accessible to collaborators, e.g., shared display [18, 101, 119]), and public 0 scale (i.e., interaction component can be seen and interacted with by arbitrary bystanders).
@@
With this dimension, we describe the autonomy of each interaction component within the overall device ecology. We coded this as flexible, semi-fixed, and fixed:
Flexible 15
device dependency indicates that all interaction components provide basic features—a single interaction component could suffice to interact with a system in a meaningful way. This can be helpful for workflows that can be easily divided properly into specific subtasks for different environments, such as data analysis workflows [66], note taking [153], or sketching [3]. The main value of HUIs in this situation comes from the interaction between the components, such as the seamless transition between components.
Semi-Fixed 40
device dependency represents systems in which one interaction component is completely independent, while others provide supplementary functionality and are thus reliant on the “main” component. This dynamic has been used primarily to extend the capabilities of existing 2D components, such as increasing the available screen space [70, 47, 87] or offering complementary views on existing content [8, 119, 54]. In contrast, a semi-fixed dependency can also be used to provide complementary capabilities to MR environments, such as providing a shared public space [76] or extracting content from 2D interaction components [131].
Fixed 27
device dependency describes systems that can only be used meaningfully with all relevant components present. With fixed device dependencies, responsibilities are exclusively distributed between components, such as using the 2D interaction component as a haptic surface for touch [18, 27, 14], for contextual interaction within the MR environment [67, 46, 126] or as a spatially aware controller [85, 96].
@@
The space dimension describes whether interaction components are co-located 83 (i.e., within the same physical space) or remote 0. Since it may be difficult to achieve a synchronous MR-HUI with remote interaction components, all the records we surveyed were exclusively co-located. An asynchronous MR-HUI with remote interaction components may be feasible but would likely forfeit potential benefits gained from any combination of complementary devices (cf. [16]). Instead, we see the potential of such remote combinations in collaborative scenarios with multiple user having individual interaction components relationships. Potential remote collaboration scenarios might include individual MR-HUIs per location and user (e.g., a HWD to visualize data in AR and a handheld device for precise input): Here, the AR visualization could be shared and synced, while the handheld devices allow individual manipulations. However, such scenarios might also require user representations (e.g., to create awareness), a communication channel, and a merge policy for conflicting input. We also discuss opportunities for collaboration using MR-HUIs in Section 6.6.
@@
Since systems in our corpus are intentionally spread across multiple interaction components, the interaction dynamics describe how each interaction component can interact with a system. Building on the BISHARE [158] design space, which classified interaction concepts as either HWD-centric or phone-centric, we identified three kinds of dynamics:
Unidirectional (2D-centric) 35
@@
Bidirectional 44
dynamics indicate that all interaction components can interact with the system equally, for example, by switching between 2D and 3D sketching [34], visualizations [103], or transferring content [133, 153].
@@
This dimension describes where the content is placed in the MR environment. We extended the design space by Reichherzer et al. [114], which categorizes content as either world-fixed or device-fixed. In addition, previous taxonomies have explored anchoring in more detail (e.g., semantic and spatial coupling in world-fixed content [37], content layout of device-fixed content [112, 23]), which we consider out of scope for this work. Our corpus is split almost equally between three general anchoring techniques:
Component-coupled 26
anchoring relates content within the MR environment to the 2D interaction components (device-fixed [114]). To create the illusion of spatial awareness and proper alignment of 2D interaction components, component-coupled usually involves spatial calibration of the 2D interaction component for stationary devices or active tracking for mobile platforms. Augmented displays and VESADs always involve a component-coupled anchor, while other systems require knowledge about the 2D interaction component for transferring content [131].
Free 30
anchoring describes MR content that is independently placed in the world or attached to real objects in the environment (world-fixed [114]). This method can be used when the 2D interaction component is not directly related to the MR environment, for example, when providing a menu in the 2D interaction component [34, 144] or a simplified but detached view of the MR content [14, 103].
Dynamic 27
anchoring may support both component-coupled and free anchoring. Previous work has explored this approach in terms of transferring content from a 2D interaction component to the MR environment or vice versa [158, 67, 153] or for cutting through 3D models [85, 96].
@@ 4.2 Emergent Trends and Opportunities
To discover possible usage patterns of different configurations, we look at the distribution and trends of dimensions 2–8 across the configuration dimension (see Fig. 5). Although existing literature covers a wide range of possibilities, we can identify several trends: (1) Spatial distribution configurations (including augmented display and VESADs) have specific requirements, making them unsuitable for serial and exclusive temporal usage as well as flexible dynamics usage. Their focus on extending a 2D interaction component is indicated by the lack of systems that demonstrate unidirectional (MR-centric) interaction dynamics and free anchoring. (2) Migratory configurations can be used regardless of their temporal dimension, but have only been explored so far within single-user systems. In this configuration, all interaction components appear to be equally important for interaction, since almost every system in our corpus uses a bidirectional interaction dynamics. (3) Unidirectional (MR-centric) interaction dynamics is only used in configurations that make use of the extended 3D capabilities of the MR interaction component (i.e., asymmetrical mirror, logical distribution, and migratory interface). (4) Although early work focused mainly on exploring complementary combinations in general logical distributions, the increase in hardware sophistication is reflected in an increase in configuration diversity: For example, symmetric configurations only appeared around 2016, while remote control configurations appeared around 2020. Since HUIs address an ever-shifting window of opportunity of contemporary hardware capabilities, the choice of possibilities is determined by the available hardware. We expect that new hardware will lead to novel combinations, creating new HUI configurations.
@@
We can identify several research opportunities by looking at gaps in current usage: (1) MR-HUIs have been exclusively explored in co-located spaces. Although this can be partially attributed to the focus on single-user systems, several systems already demonstrate the potential of MR-HUIs for collaboration. Similarly, we can see a lack of systems within public range. In both cases, future research could explore the distinct roles of each interaction component in these settings (e.g., territoriality, establishing shared and private spaces). (2) Only few systems have explored how to interact with 2D interaction components using the MR interaction component (MR-centric interaction dynamic). We see the potential in either public scenarios (e.g., avoiding hygienic issues) or dynamically enabling remote interaction with 2D interaction components [63]: Here, a symmetric configuration may be useful to offer complementary interaction possibilities. (3) We attribute the lack of systems with near range to current MR hardware limitations. For example, the limited field of view makes it hard to augment smartwatches. However, near devices could provide a complementary interface to MR HWDs by providing at-a-glance information (cf. [12]) or touch interaction. (4) Since the logical distribution configuration represents a substantial amount of records in our survey, we see potential to further differentiate this configuration. Although such a fine-grained analysis exceeds the scope of this work, future MR-HUIs might help to reveal additional patterns.
@@ 4.3 Edge Cases
Collaboration with asymmetric devices.  One of our exclusion criteria (see Sec. 3.2) was a collaboration with asymmetric devices (i.e., applications where users can interact only with one device in total). However, such systems can be situated within the cross-device taxonomy [16] and even be classified as “hybrid”, with one user on a tablet and one in AR [105, 94, 52]. We believe that many of our key characteristics are not applicable to such systems. Although collaborative systems with asymmetric devices might still provide valuable insights for MR-HUIs, we argue that this would dilute our current focus on cross-device (as opposed to “cross-user”) interaction. Instead, we see them as part of the research field of cross-reality environments (see Sec. 2.2).
@@
Reproducing reality & virtual reality.  To avoid limiting our survey to currently available technologies, we also considered systems that simulate or reproduce reality, such as VST HWDs. Such device combinations are commonly used to overcome the limitations of current OST AR HWDs (e.g., increased field of view [104], avoiding different focal planes [55, 151]), evaluate systems in scenarios that are difficult to reproduce in reality [77] (e.g., supermarkets [36]), or even simulate hardware capabilities that were not feasible at the time of publication (e.g., (transparent) tablets [85, 141]). Although this can negate some of the benefits of HUIs (e.g., high-resolution displays of current mobile devices are limited by the clarity of VST HWDs), their conceptual application design contains valuable insights for the complementary use of interaction components, regardless of the technology used. By extension, we consider combinations that employ a VR environment if they demonstrate complementary use of interaction components according to our attributes, such as using a tablet in VR [139, 34]. In contrast, we excluded papers that did not fulfill our attributes, such as ones that do not establish a mutual dependency between the virtual environment and the simulated device [36].
@@
Tangible interaction without visual output.  Mobile devices combined with AR HWDs have great potential for tangible interaction techniques that do not necessarily rely on the device’s visual output [137]. This can be useful to extend the interaction design space. However, as they do not fulfill our attributes (i.e., standalone 2D interaction component), we excluded records that do not rely on the mobile device’s screen. We see such systems in another subset of HUIs, at the intersection between the broader field of HUIs and tangible interaction [75], which map digital interactions to dedicated real-world objects.
5 How Are Mixed Reality Hybrid User Interfaces Used?
@@
In this section, we take a closer look at how MR-HUIs are used throughout our literature corpus. We highlight findings of our corpus in terms of previously used terminology (Sec. 5.1), use cases (Sec. 5.2), devices and combinations (Sec. 5.3), contribution types (Sec. 5.4), and evaluation strategies (Sec. 5.5). Finally, we discuss our survey results and provide a summary of insights (Sec. 5.6).
@@ 5.1 Terminology
We extracted the terminology used by each publication to better understand the terms used to describe MR-HUIs. Overall, we found that the author’s terminology touched the following areas: Hybrid User Interface 22, referring to the term introduced by Feiner and Shamash [47]; Hybrid ⟨other⟩ 16, referring to areas such as hybrid computing environment, hybrid display system, or hybrid setup; Cross-Device 17 and Cross-Reality Interaction 8, relating to the adjacent research areas; Multi-Device Interaction 10, describing a setup related to cross-device interaction; Augmented Displays 4, predominately used by Reipschläger et al. [118]; Transitional User Interface 2, used in relation to cross-reality collaboration or HUIs. In addition, several records used a variety of other terms 8, such as compound environment, display combination, or multimodal interaction. Lastly, the rest had no explicit terminology 11.
Figure 6: (Left) Distribution of used device types for each year within our literature corpus (years without records were omitted). Here, the complete table is shaded based on the maximum value (i.e., 9) found in it. (Right) Amount of device combinations within our corpus across a selected set of MR and 2D interaction components. Each row is shaded to show the frequency of 2D and MR component combinations

### v2

@@ 4.1 Taxonomy Dimensions
The relationship category denotes the relation between users within one or multiple systems. The vast majority of our corpus describes single-user systems 103: One user interacts with multiple complementary IO components (see Fig. 4 1 user). However, we identified several collaborative systems. We abstracted these into multi-user with individual IO components 9: Each collaborator has their own set of IO components [schroder2023collaborating, rekimoto1999augmented] (see Fig. 4 n-indiv.); and multi-user with shared IO component 14: A display is shared between collaborators as a public space [mueller2015capture, reipschlager2021personal, krug2022clear, butscher2018clusters] (see Fig. 4 n-shared).
@@
We examined the range of interaction across components as defined in the cross-device taxonomy [brudy2019crossdevice]. Since the MR IO component is commonly worn on the user’s head (e.g., AR HWD) or close to their body (e.g., handheld AR), we use this dimension to describe the scale of the 2D IO component relative to the user. In addition, the relationship can be an indicator of the range between IO components (e.g., multi-user can indicate social or public range). We differentiate between near 8 (i.e., IO component is close to the user’s body, e.g., smartwatch [grubert2015multifi, wang2014coordinated, estalagem2024wearable]), personal 92 (i.e., IO component is in the personal space, e.g., smartphone or tablet [satkowski2021insitu, hubenschmid2021stream, langner2021marvis, krug2022clear]), social 31 (i.e., IO component is in a social space accessible to collaborators, e.g., shared display [butscher2018clusters, mueller2015capture, reipschlager2021personal]), and public 2 scale (i.e., IO component can be seen and interacted with by arbitrary bystanders).
@@
With this dimension, we describe the autonomy of each IO component within the overall device ecology. We coded this as flexible, semi-fixed, and fixed:
Flexible 25
device dependency indicates that all IO components provide basic features—a single IO component could suffice to interact with a system in a meaningful way. This can be helpful for workflows that can be easily divided properly into specific subtasks for different environments, such as data analysis workflows [hubenschmid2022relive], note taking [wu2020megereality], or sketching [arora2018symbiosissketch]. The main value of HUIs in this situation comes from the interaction between the components, such as the seamless transition between components.
Semi-Fixed 52
device dependency represents systems in which one IO component is completely independent, while others provide supplementary functionality and are thus reliant on the “main” component. This dynamic has been used primarily to extend the capabilities of existing 2D components, such as increasing the available screen space [hubenschmid2023smartphone, feiner1991hybrid, langner2021marvis] or offering complementary views on existing content [berns2019myr, reipschlager2021personal, grubert2015multifi]. In contrast, a semi-fixed dependency can also be used to provide complementary capabilities to MR environments, such as providing a shared public space [james2023evaluating] or extracting content from 2D IO components [schwajda2023transforming].
Fixed 41
device dependency describes systems that can only be used meaningfully with all relevant components present. Here, responsibilities are exclusively distributed between components, such as using the 2D IO component as a haptic surface for touch [butscher2018clusters, chulpongsatorn2023holotouch, bornik2006interactive], for contextual interaction within the MR environment [hubenschmid2021stream, feiner1997touring, satkowski2022investigating] or as a spatially-aware controller [krug2022clear, luo2021exploring].
@@
The space dimension describes whether IO components are co-located 116 (i.e., within the same physical space) or remote 0. Since it may be difficult to achieve a synchronous MR-HUI with remote IO components, all the records we surveyed were exclusively co-located. An asynchronous MR-HUI with remote IO components may be feasible but would likely forfeit potential benefits gained from any combination of complementary devices (cf. [brudy2019crossdevice]). Instead, we see the potential of such remote combinations in collaborative scenarios with multiple user having individual IO components relationships. Potential remote collaboration scenarios might include individual MR-HUIs per location and user (e.g., a HWD to visualize data in AR and a handheld device for precise input): Here, the AR visualization could be shared and synced, while the handheld devices allow individual manipulations. However, such scenarios might also require user representations (e.g., to create awareness), a communication channel, and a merge policy for conflicting input. We also discuss opportunities for collaboration using MR-HUIs in Section 6.6.
@@
Since systems in our corpus are intentionally spread across multiple IO components, the interaction dynamics describe how each IO component can interact with a system. Building on the BISHARE [zhu2020bishare] design space, which classified interaction concepts as either HWD-centric or phone-centric, we identified three kinds of dynamics:
Unidirectional (2D-centric) 51
@@
Bidirectional 60
dynamics indicate that all IO components can interact with the system equally, for example, by switching between 2D and 3D sketching [drey2020vrsketchin], visualizations [nebeling2020mrat], or transferring content [serrano2015gluey, wu2020megereality].
Furthermore, we tagged every IO component by interaction paradigms that they showcase. For this, we used the non-orthogonal tags: symbolic (WIMP) [wobbrock2005guess-symbolic], reality-based-interaction (RBI) [jacob2008rbi], direct (pointing at or touching objects to interact), and indirect (interaction target is offset from the location where the input is sensed) [pfeuffer2015direct-indirect]. With a few exceptions (e.g., fixed study conditions, software toolkits), tracked HWDs and touch-enabled surfaces were tagged with RBI (85 and 59), e.g., because of pan/zoom through head/body movement and touch. We consider mid-air gestures (e.g., grab), pointing, touch, and gaze interaction as direct interaction, while the use of controllers with physical buttons, voice commands, and sign gestures (e.g., thumbs-up) are indirect interaction. Naturally, laptops and desktop computers almost exclusively relied on WIMP and indirect interaction through mouse/pointers and trackpads (42 in both cases), with a few exceptional peripheral devices. This analysis, however, did not reveal unexpected insights about, for example, underutilized IO component capabilities. Regardless, we forward the curious reader to our supplementary material for all details.
@@
This dimension describes where the content is placed in the MR environment. We extended the design space by Reichherzer et al. [reichherzer2021secondsight], which categorizes content as either world-fixed or device-fixed. In addition, previous taxonomies have explored anchoring in more detail (e.g., semantic and spatial coupling in world-fixed content [ellenberg2023spatiality], content layout of device-fixed content [rashid2012factors, cauchard2011visual]), which we consider out of scope for this work. Our corpus is split almost equally between three general anchoring techniques:
Component-coupled 32
anchoring relates content within the MR environment to the 2D IO components (device-fixed [reichherzer2021secondsight]). To create the illusion of spatial awareness and proper alignment of 2D IO components, component-coupled usually involves spatial calibration of the 2D IO component for stationary devices or active tracking for mobile platforms. Augmented displays and VESADs always involve a component-coupled anchor, while other systems require knowledge about the 2D IO component for transferring content [schwajda2023transforming].
Free 52
anchoring describes MR content that is independently placed in the world or attached to real objects in the environment (world-fixed [reichherzer2021secondsight]). This method can be used when the 2D IO component is not directly related to the MR environment, for example, when providing a menu in the 2D IO component [drey2020vrsketchin, vock2021idiar] or a simplified but detached view of the MR content [bornik2006interactive, nebeling2020mrat].
Dynamic 34
anchoring may support both component-coupled and free anchoring. Previous work has explored this approach in terms of transferring content from a 2D IO component to the MR environment or vice versa [zhu2020bishare, hubenschmid2021stream, wu2020megereality] or for cutting through 3D models [krug2022clear, luo2021exploring].
@@ 4.2 Emergent Trends and Opportunities
To discover possible usage patterns of different configurations, we look at the distribution and trends of dimensions 2–8 across the configuration dimension (see Fig. 5). Although existing literature covers a wide range of possibilities, we can identify several trends: (1) Spatial distribution configurations (including augmented display and VESADs) have specific requirements, making them unsuitable for serial and exclusive temporal usage as well as flexible dynamics usage. Their focus on extending a 2D IO component is indicated by the lack of systems that demonstrate unidirectional (MR-centric) interaction dynamics and free anchoring. (2) Migratory configurations can be used regardless of their temporal dimension, but have only been explored so far within single-user systems. In this configuration, all IO components appear to be equally important for interaction, since almost every system in our corpus uses a bidirectional interaction dynamics. (3) Unidirectional (MR-centric) interaction dynamics is only used in configurations that make use of the extended 3D capabilities of the MR IO component (i.e., asymmetrical mirror, logical distribution, and migratory interface). (4) Although early work focused mainly on exploring complementary combinations in general logical distributions, the increase in hardware sophistication is reflected in an increase in configuration diversity: For example, symmetric configurations only appeared around 2016, while remote control configurations appeared around 2020. Since HUIs address an ever-shifting window of opportunity of contemporary hardware capabilities, the choice of possibilities is determined by the available hardware. We expect that new hardware will lead to novel combinations, creating new HUI configurations.
@@
We can identify several research opportunities by looking at gaps in current usage: (1) MR-HUIs have been exclusively explored in co-located spaces. Although this can be partially attributed to the focus on single-user systems, several systems already demonstrate the potential of MR-HUIs for collaboration. Similarly, we can see a lack of systems within public range. In both cases, future research could explore the distinct roles of each IO component in these settings (e.g., territoriality, establishing shared and private spaces). (2) Only few systems have explored how to interact with 2D IO components using the MR IO component (MR-centric interaction dynamic). We see the potential in either public scenarios (e.g., avoiding hygienic issues) or dynamically enabling remote interaction with 2D IO components [horak2018when]: Here, a symmetric configuration may be useful to offer complementary interaction possibilities. (3) We attribute the lack of systems with near range to current MR hardware limitations. For example, the limited field of view makes it hard to augment smartwatches. However, near devices could provide a complementary interface to MR HWDs by providing at-a-glance information (cf. [blascheck2019glanceable]) or touch interaction. (4) Since the logical distribution configuration represents a substantial amount of records in our survey, we see potential to further differentiate this configuration. Although such a fine-grained analysis exceeds the scope of this work, future MR-HUIs might help to reveal additional patterns.
@@ 4.3 Edge Cases
Collaboration with asymmetric devices.  One of our exclusion criteria (see Sec. 3.2) was a collaboration with asymmetric devices (i.e., applications where users can interact only with one device in total). However, such systems can be situated within the cross-device taxonomy [brudy2019crossdevice] and even be classified as “hybrid”, with one user on a tablet and one in AR [olin2021designing, lunding2022exposar, gottsacker2023hybrid]. We believe that many of our key characteristics are not applicable to such systems. Although collaborative systems with asymmetric devices might still provide valuable insights for MR-HUIs, we argue that this would dilute our current focus on cross-device (as opposed to “cross-user”) interaction. Instead, we see them as part of the research field of cross-reality environments (see Sec. 2.2).
@@
Reproducing reality & virtual reality.  To avoid limiting our survey to currently available technologies, we also considered systems that simulate or reproduce reality, such as VST HWDs. Such device combinations are commonly used to overcome the limitations of current OST AR HWDs (e.g., increased field of view [normand2018enlarging], avoiding different focal planes [grubert2023text, wieland2024push2ar]), evaluate systems in scenarios that are difficult to reproduce in reality [jetter2020vr] (e.g., supermarkets [eichhorn2023shopping]), or even simulate hardware capabilities that were not feasible at the time of publication (e.g., (transparent) tablets [krug2022clear, szalavari1997personal]). Although this can negate some of the benefits of HUIs (e.g., high-resolution displays of current mobile devices are limited by the clarity of VST HWDs), their conceptual application design contains valuable insights for the complementary use of IO components, regardless of the technology used. By extension, we consider combinations that employ a VR environment if they demonstrate complementary use of IO components according to our attributes, such as using a tablet in VR [surale2019tabletinvr, drey2020vrsketchin]. In contrast, we excluded papers that did not fulfill our attributes, such as ones that do not establish a mutual dependency between the virtual environment and the simulated device [eichhorn2023shopping].
@@
Tangible interaction without visual output.  Mobile devices combined with AR HWDs have potential for tangible interaction techniques that do not necessarily rely on the device’s visual output [stellmacher2024exploring]. This can be useful to extend the interaction design space. However, as they do not fulfill our attributes (i.e., standalone 2D IO component), we excluded records that do not rely on the mobile device’s screen. We see such systems in another subset of HUIs, at the intersection of the broader field of HUIs and tangible interaction [ishii2008tangible], which map digital interactions to real-world objects.
5 How Are Mixed Reality HUIs Used?
@@
In this section, we take a closer look at how MR-HUIs are used throughout our literature corpus. We highlight findings of our corpus in terms of previously used terminology (Sec. 5.1), use cases (Sec. 5.2), devices and combinations (Sec. 5.3), contribution types (Sec. 5.4), and evaluation strategies (Sec. 5.5). Finally, we provide a summary of insights (Sec. 5.6) and discuss the potential and pitfalls of HUIs (Sec. 5.7).
@@ 5.1 Terminology
We extracted the terminology used by each publication to better understand the terms used to describe MR-HUIs. Overall, we found that the author’s terminology touched the following areas: Hybrid User Interface 29, referring to the term introduced by Feiner and Shamash [feiner1991hybrid]; Hybrid ⟨other⟩ 20, referring to areas such as hybrid computing environment, hybrid display system, or hybrid setup; Cross-Device 26 and Cross-Reality Interaction 14, relating to the adjacent research areas; Multi-Device Interaction 10, describing a setup related to cross-device interaction; Augmented Displays 4, predominately used by Reipschläger et al. [reipschlager2020augmented]; Transitional User Interface 4, used in relation to cross-reality collaboration or HUIs. In addition, several records used a variety of other terms 13, such as compound environment, display combination, or multimodal interaction. Lastly, the rest had no explicit terminology 20.
Figure 6: (Left) Distribution of used device types for each year within our literature corpus (years without records were omitted). Here, the complete table is shaded based on the maximum value (i.e., 10) found in it. (Right) Amount of device combinations within our corpus across a selected set of MR and 2D IO components. Each row is shaded to show the frequency of 2D and MR component combinations.